# Med3DVLM: An Efficient Vision-Language Model for 3D Medical Image Analysis

### v2

@@ 3.3 Evaluation Metrics
We evaluate the quality of radiology report generation and open-ended VQA using four key metrics: BLEU [25], ROUGE [17], METEOR [3], and BERTScore [46]. BLEU measures precision by evaluating the overlap of n-grams between the generated text and the reference text, focusing on how many words or phrases in the generated report are found in the reference text. The BLEU score is calculated as:
@@ 3.4 Implementation Details
We implemented Med3DVLM using PyTorch [26] and Hugging Face Transformers [38], leveraging DeepSpeed ZeRO2 and BF16 precision on 8 NVIDIA A100 80GB GPUs. We used the AdamW optimizer [21] with a warmup ratio of 0.03 and a cosine learning rate scheduler.
@@ 3.4.1 Contrastive Pretraining
During contrastive pretraining on the M3D-Cap dataset, we used image-text pairs, where images were 3D CT volumes and texts were corresponding radiology reports. We computed similarity scores between pooled image and text embeddings to optimize a contrastive loss. We resized 3D CT volumes to $128\times 256\times 256$. We used DCFormer-small [1] as the vision encoder, and ClinicalBERT [34] as the text encoder. Image features were aggregated via mean pooling, while text features were achieved from the [CLS] token, as it captures the global semantic representation of the entire sequence in ClinicalBERT, making it suitable for contrastive alignment with image features. Both were projected to a shared 768-dimensional space. We used a batch size of 64, learning rate of $1\times 10^{-4}$, weight decay of 0.1, and trained for 100 epochs.
@@ 3.4.3 VLM Fine-tuning
For VLM fine-tuning, We continued training using image–question–answer triplets from M3D-Cap and M3D-VQA, now including yes/no questions. The image encoder and MLP-Mixer projector were reused. The model was fine-tuned to generate answers conditioned on the image and question. We used LoRA [13] with a rank of 16, scaling factor $\alpha=32$, and dropout of 0.05, updating only the LoRA modules and the projector while keeping all other weights frozen. We fine-tuned the model using a batch size of 8, learning rate of $5\times 10^{-5}$, no weight decay, and training for 5 epochs. All M3D-LaMed results are reported as in the original paper [2].

### v3

@@ 3.3 Evaluation Metrics
We evaluate the quality of radiology report generation and open-ended VQA using four key metrics: BLEU [papineni2002bleu], ROUGE [lin2004rouge], METEOR [banerjee2005meteor], and BERTScore [BERT-Score]. BLEU measures precision by evaluating the overlap of n-grams between the generated text and the reference text, focusing on how many words or phrases in the generated report are found in the reference text. The BLEU score is calculated as:
@@ 3.4 Implementation Details
We implemented Med3DVLM using PyTorch [paszke2019pytorch] and Hugging Face Transformers [wolf2019huggingface], leveraging DeepSpeed ZeRO2 and BF16 precision on 8 NVIDIA A100 80GB GPUs. We used the AdamW optimizer [loshchilov2017decoupled] with a warmup ratio of 0.03 and a cosine learning rate scheduler.
@@ 3.4.1 Contrastive Pretraining
During contrastive pretraining on the M3D-Cap dataset, we used image-text pairs, where images were 3D CT volumes and texts were corresponding radiology reports. We computed similarity scores between pooled image and text embeddings to optimize a contrastive loss. We resized 3D CT volumes to $128\times 256\times 256$. We used DCFormer-small [ates2025dcformer] as the vision encoder, and ClinicalBERT [wang2023optimized] as the text encoder. Image features were aggregated via mean pooling, while text features were achieved from the [CLS] token, as it captures the global semantic representation of the entire sequence in ClinicalBERT, making it suitable for contrastive alignment with image features. Both were projected to a shared 768-dimensional space. We used a batch size of 64, learning rate of $1\times 10^{-4}$, weight decay of 0.1, and trained for 100 epochs.
@@ 3.4.3 VLM Fine-tuning
For VLM fine-tuning, We continued training using image–question–answer triplets from M3D-Cap and M3D-VQA, now including yes/no questions. The image encoder and MLP-Mixer projector were reused. The model was fine-tuned to generate answers conditioned on the image and question. We used LoRA [hu2022lora] with a rank of 16, scaling factor $\alpha=32$, and dropout of 0.05, updating only the LoRA modules and the projector while keeping all other weights frozen. We fine-tuned the model using a batch size of 8, learning rate of $5\times 10^{-5}$, no weight decay, and training for 5 epochs. All M3D-LaMed results are reported as in the original paper [bai2024m3d].